# Total Failure of Approachability at Successors of Singulars of Countable Cofinality

### v1

@@ 2. Iteration Theorems for Prikry-Type forcings
Let $(\mathbb{P},\leq)$ be a partial order and let $\leq_{0}$ be another partial order on $\mathbb{P}$ such that $\leq$ refines $\leq_{0}$. We say that $(\mathbb{P},\leq,\leq_{0})$ is a Prikry-type forcing if for every $p\in\mathbb{P}$ and every statement $\sigma$ in the forcing language there is $q\leq_{0}p$ which decides $\sigma$ (this is known as the Prikry property).
@@
Other examples include variants of Namba forcing. The original Namba forcing (see [BCH90]) shares some similarities with Prikry forcing. It adds a countable cofinal sequence to $\aleph_{2}$ without collapsing $\aleph_{1}$ and without adding reals (assuming $\mathsf{CH}$ holds in the ground model). However, in the case of Namba forcing, both of these regularity properties follow from abstract considerations (such as Shelah’s $\mathbb{I}$-condition, see [She17, Chapter XI]) which introduce complications regarding the preservation of larger cardinals and iterability. To remedy both of these problems, we will be using Namba forcings defined from particularly well-behaved ideals which function more similarly to supercompact Prikry forcing.
We start with the problem of iterating Prikry-type forcings. This was largely solved by Magidor in his famous paper proving that it is consistent that the first strongly compact cardinal can also be the first measurable cardinal (which he achieved by successively singularizing every smaller measurable cardinal; see [Mag76]). We will be using iterations both with full support and with Easton support. The most important point that makes sure the iterations have the Prikry property is the restriction on where non-direct extensions can be used: For full support, only finitely many non-direct extensions are allowed while for Easton support, arbitrarily many non-direct extensions are allowed outside the previous support and only finitely many non-direct extensions are allowed inside the previous support.
@@
Let $((\mathbb{P}_{\alpha},\leq_{\alpha},\leq_{\alpha,0}),(\dot{\mathbb{Q}}_{\alpha},\dot{\leq}_{\alpha},\dot{\leq}_{\alpha,0}))_{\alpha<\rho}$ be a sequence such that each $(\mathbb{P}_{\alpha},\leq_{\alpha},\leq_{\alpha,0})$ is a poset and each $(\dot{\mathbb{Q}}_{\alpha},\dot{\leq}_{\alpha},\dot{\leq}_{\alpha,0})$ is a $\mathbb{P}_{\alpha}$-name for a Prikry-type poset. We define the statement $((\mathbb{P}_{\alpha},\leq_{\alpha},\leq_{\alpha,0}),(\dot{\mathbb{Q}}_{\alpha},\dot{\leq}_{\alpha},\dot{\leq}_{\alpha,0}))_{\alpha<\rho}$ is a full support Magidor iteration of Prikry-type forcings of length $\rho$” by induction on $\rho$.
@@
$p\upharpoonright\rho^{\prime}\Vdash\sigma\in[\check{\mu}]^{<\check{\delta}}\wedge\dot{q}(\rho^{\prime})\Vdash\tau\subseteq\check{\sigma}$
@@
$\sigma_{\alpha}:=\exists r\in\dot{\mathbb{P}}_{\alpha,\rho}\exists x\in[\check{\mu}]^{<\check{\delta}}(r\leq_{(\alpha,\rho),0}p\upharpoonright[\alpha,\rho)\wedge r\Vdash\tau\subseteq\check{x})$
@@
$q\upharpoonright\alpha\Vdash q(\alpha)\Vdash\dot{r}\Vdash\tau\subseteq\check{x}$
@@
In general, working with the direct extension ordering on a Magidor iteration is quite difficult since it does not quite behave like an iteration. Due to this, we will be using the following idea: Whenever $(\mathbb{P},\leq,\leq_{0})$ is a Prikry-type poset and $(\dot{\mathbb{Q}},\dot{\leq},\dot{\leq}_{0})$ is a $\mathbb{P}$-name for a poset, $(\dot{\mathbb{Q}},\dot{\leq}_{0})$ is a $(\mathbb{P},\leq)$-name for a poset (since the ordering $\leq_{0}$ on $\mathbb{P}$ is just used to derive some regularity properties). As such, by Lemma 1.19, the identity function is a projection from $(\mathbb{P},\leq)\times\mathbb{T}((\mathbb{P},\leq),(\dot{\mathbb{Q}},\dot{\leq}_{0}))$ onto $(\mathbb{P},\leq)*(\dot{\mathbb{Q}},\dot{\leq}_{0})$. However, since $\leq$ refines $\leq_{0}$, we also obtain a projection from $(\mathbb{P},\leq_{0})\times(\mathbb{T}((\mathbb{P},\leq),(\dot{\mathbb{Q}},\dot{\leq}_{0})))$ onto the direct extension ordering on $\mathbb{P}*\dot{\mathbb{Q}}$:
@@
Now we show that the identity is a projection. Let $p\in\mathbb{P}$ and $t\leq_{\mathbb{P}}p$. We want to find $p^{\prime}\leq_{\mathbb{T}}p$ with $p^{\prime}\leq_{\mathbb{P}}t$. By induction on $\alpha$, let $p^{\prime}(\alpha)$ be a $(\mathbb{P}_{\alpha},\leq_{\alpha})$-name for an element of $\dot{\mathbb{Q}}_{\alpha}$ forced (over $(\mathbb{P}_{\alpha},\leq_{\alpha})$) by $p^{\prime}\upharpoonright\alpha$ to be equal to $t(\alpha)$ and by conditions incompatible with $p^{\prime}\upharpoonright\alpha$ to be equal to $p(\alpha)$. Then clearly $p^{\prime}\leq_{\mathbb{T}}p$, since $p^{\prime}(\alpha)$ is always forced to be below $p(\alpha)$ and $p^{\prime}\leq_{0}t$: By induction, $p^{\prime}\upharpoonright\alpha\leq 0t\upharpoonright\alpha$. Since $t\upharpoonright\alpha\Vdash t(\alpha)\leq_{0}p(\alpha)$, $p^{\prime}\upharpoonright\alpha$ forces $p^{\prime}(\alpha)=t(\alpha)\leq_{0}p(\alpha)$. It follows that $p^{\prime}\leq_{0}pt$. ∎
@@ 3. The Laver-Ideal Property
Classically, by work of Laver (unpublished) and Galvin-Jech-Magidor (see [GJM78, Theorem 2]), whenever $\mu$ is regular and $\nu>\mu$, the Levy collapse $\operatorname{Coll}(\mu,<\nu)$, which forces $\nu=\mu^{+}$, also forces $\mathsf{LIP}(\mu,\nu,\nu)$. In this section, we prove a characterization for posets which force $\mathsf{LIP}$ (which basically follows from their proof, combined with ideas of Foreman) and show that when starting from a supercompact cardinal instead of a measurable cardinal, we even obtain $\mathsf{LIP}(\mu,\nu,\kappa)$ for all $\kappa\geq\nu$. We do this because we later have to show that the direct extension ordering on the tail of our iteration forces enough instances of $\mathsf{LIP}$ to make other parts of the iteration work (this will become apparent in the proof of the main theorem).
@@
Let $(X_{i})_{i<\gamma}$ ($\gamma<\nu$) be elements of $I$. Let $H$ be any $j(\mathbb{P})/G$-generic filter over $V[G]$. Then $j^{+}((X_{i})_{i<\gamma})=(j^{+}(X_{i}))_{i<\gamma}$ and so
@@
i.e. $\bigcup_{i<\gamma}X_{i}\in I$. Normality follows the same way. ∎
@@
Let $G_{\alpha}$ be $\mathbb{P}\cap V_{\alpha}$-generic. Let $X\in V[G_{\alpha}]$ be a weakly directed subset of $\mathbb{P}/G_{\alpha}$ of size $\mu$. It follows that $X\in V$. It also follows that $\pi[X]\subseteq\mathbb{P}\cap V_{\alpha}$ (where $\pi\colon\mathbb{P}\to(\mathbb{P}\cap V_{\alpha})$ is the projection) and so there is $p\in G_{\alpha}$ with $p\leq\pi(q)$ for every $q\in X$. By Definition 3.2 (3), there is $p^{*}\in\mathbb{P}$ such that $\pi(p^{*})=p$ and $p\leq q$ for every $q\in X$. Ergo, $p^{*}\in\mathbb{P}/G_{\alpha}$ and $p^{*}$ is a lower bound of $X$. ∎
@@
$j[\kappa]\in j^{+}\left(\bigcap_{i<\gamma}B(p_{i},\dot{q}_{i},\tau_{i})\right)=\bigcap_{i<\gamma}j^{+}(B(p_{i},\dot{q}_{i},\tau_{i}))$
@@ 4. Collapsing without Approachability
As stated in the discussion below Lemma 1.11, the main obstacle to obtaining a model where there are many non-approachable points in the successor of a singular cardinal is finding a poset which collapses the successor of a singular cardinal of a given cofinality without making it approachable with respect to any normal, subadditive coloring. To achieve this for $\kappa=\mu^{+}$, $\mu$ singular, Shelah used $\operatorname{Coll}(\operatorname{cf}(\mu),\mu)$. This poset has size $\mu$ (under $\mathsf{GCH}$) and thus preserves $\mu^{+}$ and, in particular, does not make it approachable with respect to any normal, subadditive coloring. However, this approach has two major drawbacks: On one hand, it is specific to obtaining non-approachable points of cofinality $\operatorname{cf}(\mu)^{+}$ (since larger collapses would also collapse $\kappa$) and it cannot be iterated (since $\kappa$ always needs to be preserved). As such, it was a major open problem whether there could be a poset which can collapse $\kappa$ as a cardinal without making it approachable with respect to any normal subadditive coloring.
A positive solution to this problem was found by the author and Levine for successors of singulars of countable cofinality (see [JL25]). In the cited work, we relied on having $\mathsf{LIP}(\mu,\aleph_{n},\aleph_{n})$ for infinitely many $n\in\omega$. In this section we will show that it is possible to obtain a poset which accomplishes the same task but can be defined from weaker hypotheses (i.e. assuming $\mathsf{LIP}(\mu,\aleph_{k},\aleph_{n})$ for a single $k$ and infinitely many $n\in\omega$). This is crucial in making the later iteration work since we will just need a single generically supercompact cardinal instead of infinitely many generically measurable cardinals which is much easier to obtain in practice.
We first introduce a variant of Namba forcing which we call supercompact Namba forcing. This forcing functions similarly to supercompact Prikry forcing but can be defined for successor cardinals.
@@
    $\operatorname{osucc}_{p}(s):=\{x\in[\kappa_{|s|}]^{<\mu^{+}}\;|\;s^{\frown}x\in p\}\in B_{|s|}$
@@
By the density of each $B_{n}$, it does not matter whether we require $\operatorname{osucc}_{p}(s)\in B_{|s|}$ or $\operatorname{osucc}_{p}(s)\in I_{|s|}^{+}$ so we will be ignoring this distinction (more precisely, any tree which splits into $I_{|s|}^{+}$ can be refined to one which splits into $B_{|s|}$).
The forcing functions similarly to regular Namba forcing defined using a sequence $(I_{n})_{n\in\omega}$ of ideals on $(\kappa_{n})_{n\in\omega}$. The main advantage (which is crucial for this paper) is that we are only assuming a single instance of $\mathsf{LIP}$. Despite this, the usual results for Namba forcing still apply (we will follow [CK18, Section 6]). To simplify our presentation, we fix the following:
@@
    $\operatorname{osucc}_{p}^{n}(t):=\{r\in\prod_{i\in[|t|,|t|+n-1]}[\kappa_{i}]^{<\mu^{+}}\;|\;t^{\frown}r\in p\}$
@@
We first prove that $(\mathbb{L}(\mu,(\kappa_{n},I_{n},B_{n})_{n\in\omega}),\leq,\leq_{0})$ is a Prikry-type forcing and derives its regularity properties from a sufficiently closed direct extension ordering.
@@
Assume toward a contradiction that $n>|\operatorname{stem}(p)|$. We will find $r\leq_{0}q$ such that whenever $s\in r$, |s|=n1|s|=n-1, $r\upharpoonright s\in D$, thus contradicting the minimality of $n$. Let $s\in q$, |s|=n1|s|=n-1. Since $|s|\geq|\operatorname{stem}(p)|$, $\operatorname{osucc}_{q}(s)\in I_{|s|}^{+}$. For any $x\in\operatorname{osucc}_{q}(s)$, there is $\gamma_{x}$ such that $q\upharpoonright(s^{\frown}x)\Vdash\tau=\check{\gamma}_{x}$. By the $\mu^{+}-$completeness of $I_{|s|}^{+}$, there is $B_{s}\subseteq\operatorname{osucc}_{q}(s)$ and $\gamma_{s}$ such that for any $\alpha\in B_{s}$, $q\upharpoonright(s^{\frown}x)\Vdash\tau=\check{\gamma}_{s}$. Now let $r^{s}:=\{t\in q\upharpoonright s\;|\;|t|\leq|s|\vee t(|s|)\in B_{s}\}$. It follows that $r^{s}\leq_{0}q\upharpoonright(s^{\frown}\alpha)$ and $r^{s}\Vdash\tau=\check{\gamma}_{s}$, i.e. $r^{s}\in D$. Let $r:=\bigcup_{s\in q,|s|=n-1}r^{s}$. Then $r\leq_{0}q$ and whenever $s\in r$, |s|=n1|s|=n-1, $r\upharpoonright s=r^{s}$ and thus $r\upharpoonright s\in D$. This $r$ is what we wanted to construct. ∎
As is common with Prikry-type forcings, $\mathbb{L}(\mu,(\kappa_{n},I_{n},B_{n})_{n\in\omega})$ derives its regularity properties from the Prikry property together with a sufficiently closed direct extension ordering. For technical reasons we need that the direct extension ordering is strongly $\mu-$directed closed (most importantly, we need that statement to show that our later iteration forces $\mathsf{LIP}$). To show this, we first prove some elementary statements regarding ideal combinatorics.
@@
$X\times Y\supseteq A\in I\otimes J:\Leftrightarrow\{x\in X\;|\;\{y\in Y\;|\;(x,y)\in A\}\in J^{+}\}\in I$
@@
In particular, we can write expressions such as $I_{0}\otimes\dots\otimes I_{n}$” without having to worry about the order in which these products are taken.
The Fubini product is related to Namba forcing as follows: Given finitely many sets $(X_{k})_{k<n}$ and ideals $(J_{k})_{k<n}$, we say that $p\subseteq\bigcup_{l<n}\prod_{k\leq l}X_{k}$ is a $(J_{k})_{k<n}$-Laver-tree if $p$ is nonempty, closed under restriction and for every $s\in p$ with $|s|<n$, $\operatorname{osucc}_{p}(s):=\{x\in X_{|s|}\;|\;s^{\frown}x\in p\}\in J_{|s|}^{+}$. We let $p]:=\{s\in p\;|\;|s|=n\}$.
@@
Let $J_{0},\dots,J_{n-1}$ be ideals. Then a set $A\subseteq X_{0}\times\dots\times X_{n-1}$ is in $(J_{0}\otimes\dots\otimes J_{n-1})^{+}$ if and only if there is a $(J_{k})_{k<n}$-Laver-tree $p$ such that $p]\subseteq A$.

### v2

@@ 2. Iteration Theorems for Prikry-Type forcings
Let $(\mathbb{P},\leq)$ be a partial order and let $\leq_{0}$ be another partial order on $\mathbb{P}$ such that $\leq$ refines $\leq_{0}$. We say that $(\mathbb{P},\leq,\leq_{0})$ is a Prikry-type forcing if for every $p\in\mathbb{P}$ and every statement $\sigma$ in the forcing language there is $q\leq_{0}p$ which decides $\sigma$. The latter statement is known as the Prikry property.
@@
Other examples include variants of Namba forcing. The original Namba forcing (see [BCH90]) shares some similarities with Prikry forcing. It adds a countable cofinal sequence to $\aleph_{2}$ without collapsing $\aleph_{1}$. If $\mathsf{CH}$ holds, it is even true that no new reals are added. However, in the case of Namba forcing, both of these regularity properties follow from abstract considerations (such as Shelah’s $\mathbb{I}$-condition, see [She17, Chapter XI]) which introduce complications regarding the preservation of larger cardinals and iterability. To remedy both of these problems, we will be using Namba forcings defined from particularly well-behaved ideals which function more similarly to supercompact Prikry forcing.
We start with the problem of iterating Prikry-type forcings. This was largely solved by Magidor in his famous paper proving that it is consistent that the first strongly compact cardinal can also be the first measurable cardinal. He achieved this by successively singularizing every smaller measurable cardinal (see [Mag76]). We will be using iterations both with full support and with Easton support. The most important point that makes sure the iterations have the Prikry property is the restriction on where non-direct extensions can be used: For full support, only finitely many non-direct extensions are allowed while for Easton support, arbitrarily many non-direct extensions are allowed outside the previous support and only finitely many non-direct extensions are allowed inside the previous support.
@@
Let $((\mathbb{P}_{\alpha},\leq_{\alpha},\leq_{\alpha,0}),(\dot{\mathbb{Q}}_{\alpha},\dot{\leq}_{\alpha},\dot{\leq}_{\alpha,0}))_{\alpha<\rho}$ be a sequence such that each $(\mathbb{P}_{\alpha},\leq_{\alpha},\leq_{\alpha,0})$ is a Prikry-type poset and each $(\dot{\mathbb{Q}}_{\alpha},\dot{\leq}_{\alpha},\dot{\leq}_{\alpha,0})$ is a $\mathbb{P}_{\alpha}$-name for a Prikry-type poset. We define the statement $((\mathbb{P}_{\alpha},\leq_{\alpha},\leq_{\alpha,0}),(\dot{\mathbb{Q}}_{\alpha},\dot{\leq}_{\alpha},\dot{\leq}_{\alpha,0}))_{\alpha<\rho}$ is a full support Magidor iteration of Prikry-type forcings of length $\rho$” by induction on $\rho$.
@@
$p\upharpoonright\rho^{\prime}\Vdash\sigma\in[\check{\mu}]^{<\check{\delta}}\wedge\dot{q}(\rho^{\prime})\Vdash\tau\subseteq\check{\sigma}.$
@@
$\sigma_{\alpha}:=\exists r\in\dot{\mathbb{P}}_{\alpha,\rho}\exists x\in[\check{\mu}]^{<\check{\delta}}(r\leq_{(\alpha,\rho),0}p\upharpoonright[\alpha,\rho)\wedge r\Vdash\tau\subseteq\check{x}).$
@@
$q\upharpoonright\alpha\Vdash q(\alpha)\Vdash\dot{r}\Vdash\tau\subseteq\check{x}.$
@@
In general, working with the direct extension ordering on a Magidor iteration is quite difficult since it does not quite behave like an iteration. Due to this, we will be using the following idea: Whenever $(\mathbb{P},\leq,\leq_{0})$ is a Prikry-type poset and $(\dot{\mathbb{Q}},\dot{\leq},\dot{\leq}_{0})$ is a $\mathbb{P}$-name for a poset, $(\dot{\mathbb{Q}},\dot{\leq}_{0})$ is a $(\mathbb{P},\leq)$-name for a poset. [^4] The ordering $\leq_{0}$ on $\mathbb{P}$ is just used to derive some regularity properties As such, by Lemma 1.20, the identity function is a projection from $(\mathbb{P},\leq)\times\mathbb{T}((\mathbb{P},\leq),(\dot{\mathbb{Q}},\dot{\leq}_{0}))$ onto $(\mathbb{P},\leq)*(\dot{\mathbb{Q}},\dot{\leq}_{0})$. However, since $\leq$ refines $\leq_{0}$, we also obtain a projection from $(\mathbb{P},\leq_{0})\times(\mathbb{T}((\mathbb{P},\leq),(\dot{\mathbb{Q}},\dot{\leq}_{0})))$ onto the direct extension ordering on $\mathbb{P}*\dot{\mathbb{Q}}$:
@@
Now we show that the identity is a projection. Let $p\in\mathbb{P}$ and $t\leq_{\mathbb{P}}p$. We want to find $p^{\prime}\leq_{\mathbb{T}}p$ with $p^{\prime}\leq_{\mathbb{P}}t$. By induction on $\alpha$, let $p^{\prime}(\alpha)$ be a $(\mathbb{P}_{\alpha},\leq_{\alpha})$-name for an element of $\dot{\mathbb{Q}}_{\alpha}$ forced (over $(\mathbb{P}_{\alpha},\leq_{\alpha})$) by $p^{\prime}\upharpoonright\alpha$ to be equal to $t(\alpha)$ and by conditions incompatible with $p^{\prime}\upharpoonright\alpha$ to be equal to $p(\alpha)$. Then clearly $p^{\prime}\leq_{\mathbb{T}}p$, since $p^{\prime}(\alpha)$ is always forced to be below $p(\alpha)$ and $p^{\prime}\leq_{0}t$: By induction, $p^{\prime}\upharpoonright\alpha\leq_{0}t\upharpoonright\alpha$. Since $t\upharpoonright\alpha\Vdash t(\alpha)\leq_{0}p(\alpha)$, $p^{\prime}\upharpoonright\alpha$ forces $p^{\prime}(\alpha)=t(\alpha)\leq_{0}p(\alpha)$. It follows that $p^{\prime}\leq_{0}p$. ∎
@@ 3. The Laver-Ideal Property
Classically, by work of Laver (unpublished) and Galvin-Jech-Magidor (see [GJM78, Theorem 2]), whenever $\mu$ is regular and $\nu>\mu$, the Levy collapse $\operatorname{Coll}(\mu,<\nu)$, which forces $\nu=\mu^{+}$, also forces $\mathsf{LIP}(\mu,\nu,\nu)$. In this section, we prove a characterization for posets which force $\mathsf{LIP}$ (which basically follows from their proof, combined with ideas of Foreman) and show that when starting from a supercompact cardinal instead of a measurable cardinal, we even obtain $\mathsf{LIP}(\mu,\nu,\kappa)$ for all $\kappa\geq\nu$. As will become apparent in the proof of the main theorem, we do this because we later have to show that the direct extension ordering on the tail of our iteration forces enough instances of $\mathsf{LIP}$ to make other parts of the iteration work.
@@
Let $(X_{i})_{i<\gamma}$ ($\gamma<\nu$) be a sequence of elements of $I$. Let $H$ be any $j(\mathbb{P})/G$-generic filter over $V[G]$. Then $j^{+}((X_{i})_{i<\gamma})=(j^{+}(X_{i}))_{i<\gamma}$ and so
@@
i.e. $\bigcup_{i<\gamma}X_{i}\in I$.
For normality, let $A\in I^{+}$ and let $f$ be a regressive function on $A$. Let $H$ be a $j(\mathbb{P})/G$-generic filter over $V[G]$ such that $j[\kappa]\in j^{+}(A)$. Since $f$ is a regressive function on $A$, $j(f)$ is a regressive function on $j(A)$ and so $j(f)(j[\kappa])\in j[\kappa]$. Let $j(f)(j[\kappa])=j(\alpha)$ for some $\alpha\in\kappa$. Then, clearly,
$j[\kappa]\in j^{+}\left(\{x\in A\;|\;f(x)=\alpha\}\right)$
and thus $f$ is constant on the set $\{x\in A\;|\;f(x)=\alpha\}$ which is a member of $I^{+}$. ∎
@@
Let $G_{\alpha}$ be $\mathbb{P}\cap V_{\alpha}$-generic. Let $X\in V[G_{\alpha}]$ be a weakly directed subset of $\mathbb{P}/G_{\alpha}$ of size $\mu$. It follows that $X\in V$. It also follows that $\pi[X]\subseteq\mathbb{P}\cap V_{\alpha}$ – where $\pi\colon\mathbb{P}\to(\mathbb{P}\cap V_{\alpha})$ is the projection – and so there is $p\in G_{\alpha}$ with $p\leq\pi(q)$ for every $q\in X$. By Definition 3.2 (3), there is $p^{*}\in\mathbb{P}$ such that $\pi(p^{*})=p$ and $p\leq q$ for every $q\in X$. Ergo, $p^{*}\in\mathbb{P}/G_{\alpha}$ and $p^{*}$ is a lower bound of $X$. ∎
@@
$j[\kappa]\in j^{+}\left(\bigcap_{i<\gamma}B(p_{i},\dot{q}_{i},\tau_{i})\right)=\bigcap_{i<\gamma}j^{+}(B(p_{i},\dot{q}_{i},\tau_{i})).$
@@ 4. Collapsing without Approachability
As stated in the discussion below Lemma 1.12, the main obstacle to obtaining a model where there are many non-approachable points in the successor of a singular cardinal is finding a poset which collapses the successor of a singular cardinal of a given cofinality without making it approachable with respect to any normal, subadditive coloring. To achieve this for $\kappa=\mu^{+}$, $\mu$ singular, Shelah used $\operatorname{Coll}(\operatorname{cf}(\mu),\mu)$. Assuming the $\mathsf{GCH}$, this poset has size $\mu$ and thus preserves $\mu^{+}$. In particular, it does not make $\mu^{+}$ approachable with respect to any normal, subadditive coloring. However, this approach has two major drawbacks: On one hand, since larger collapses would also collapse $\kappa$, it is specific to obtaining non-approachable points of cofinality $\operatorname{cf}(\mu)^{+}$ and, since $\kappa$ always needs to be preserved, it cannot be iterated. As such, it was a major open problem whether there could be a poset which can collapse $\kappa$ as a cardinal without making it approachable with respect to any normal subadditive coloring.
A positive solution to this problem was found by the author and Levine for successors of singulars of countable cofinality (see [JL25]). In the cited work, we relied on having $\mathsf{LIP}(\mu,\aleph_{n},\aleph_{n})$ for infinitely many $n\in\omega$. In this section we will show that it is possible to obtain a poset which accomplishes the same task but can be defined from weaker hypotheses – assuming $\mathsf{LIP}(\mu,\aleph_{k},\aleph_{n})$ for a single $k$ and infinitely many $n\in\omega$. This is crucial in making the later iteration work since we will just need a single generically supercompact cardinal instead of infinitely many generically measurable cardinals which is much easier to obtain in practice.
We first introduce our new variant of Namba forcing. Just like in the case of supercompact Prikry forcing, our trees will consist of sequences of elements of some $\lambda]^{<\kappa}$ as opposed to sequences of ordinals. However, unlike supercompact Prikry forcing, it can be defined for successor cardinals since we are using $\mathsf{LIP}$-ideals as opposed to supercompact measures.
@@
    $\operatorname{osucc}_{p}(s):=\{x\in[\kappa_{|s|}]^{<\mu^{+}}\;|\;s^{\frown}x\in p\}\in B_{|s|}.$
@@
Since each $B_{n}$ is dense in $I_{n}^{+}$, we can refine trees which split into elements of $I_{n}^{+}$ to conditions in $\mathbb{L}(\mu,(\kappa_{n},I_{n},B_{n})_{n\in\omega})$:
Remark 4.2.
Let $p\subseteq\bigcup_{n\in\omega}\prod_{k<n}[\kappa_{k}]^{<\mu^{+}}$ be closed under restriction such that there is some $\operatorname{stem}(p)\in p$ such that for all $s\in p$, $\operatorname{stem}(p)\subseteq s$ or $s\subseteq\operatorname{stem}(p)$ and whenever $\operatorname{stem}(p)\subseteq s$,
$\operatorname{osucc}_{p}(s)\in I_{|s|}^{+}.$
For each $s\in p$ with $\operatorname{stem}(p)\subseteq s$, choose $A_{s}\subseteq\operatorname{osucc}_{p}(s)$ with $A_{s}\in B_{|s|}$. Let $q$ consist of all $t\in p$ such that whenever $s$ is an initial segment of $t$ with $\operatorname{stem}(p)\subseteq s$, $t(|s|)\in A_{s}$. It follows that, for each $t\in q$, $\operatorname{osucc}_{q}(t)=A_{t}$. Therefore, $q\subseteq p$ and $q\in\mathbb{L}(\mu,(\kappa_{n},I_{n},B_{n})_{n\in\omega})$.
Ergo, in many constructions, we will only aim to produce a tree with $(I_{n})_{n\in\omega}$-positive splitting sets, knowing that any such tree can be refined into a condition in $\mathbb{L}(\mu,(\kappa_{n},I_{n},B_{n})_{n\in\omega})$.
The forcing functions similarly to regular Namba forcing defined using a sequence $(I_{n})_{n\in\omega}$ of ideals on $(\kappa_{n})_{n\in\omega}$. The main advantage (which is crucial for this paper) is that we are only assuming a single instance of $\mathsf{LIP}$. Despite this, our Namba forcing has all of the necessary regularity properties, most importantly the strong Prikry property (see Lemma 4.4), the ability to decide names for small ordinals using direct extensions (see Lemma 4.5) and a sufficiently strongly directed-closed direct extension ordering (see 4.10). For the first two results, we will follow Cox-Krueger (see [CK18]). Previously, similar results were obtained by Cummings and Magidor (see [CM11]).
@@
    $\operatorname{osucc}_{p}^{n}(t):=\{r\in\prod_{i\in[|t|,|t|+n-1]}[\kappa_{i}]^{<\mu^{+}}\;|\;t^{\frown}r\in p\}.$
@@
We first prove that $(\mathbb{L}(\mu,(\kappa_{n},I_{n},B_{n})_{n\in\omega}),\leq,\leq_{0})$ is a Prikry-type forcing and derive its regularity properties from a sufficiently closed direct extension ordering.
@@
Assume toward a contradiction that $n>|\operatorname{stem}(p)|$. We will find $r\leq_{0}q$ such that whenever $s\in r$, |s|=n1|s|=n-1, $r\upharpoonright s\in D$, thus contradicting the minimality of $n$. Let $s\in q$, |s|=n1|s|=n-1. Since $|s|\geq|\operatorname{stem}(p)|$, $\operatorname{osucc}_{q}(s)\in I_{|s|}^{+}$. For any $x\in\operatorname{osucc}_{q}(s)$, there is $\gamma_{x}$ such that $q\upharpoonright(s^{\frown}x)\Vdash\tau=\check{\gamma}_{x}$. By the $\mu^{+}-$completeness of $I_{|s|}$, there is $B_{s}\subseteq\operatorname{osucc}_{q}(s)$ and $\gamma_{s}$ such that for any $\alpha\in B_{s}$, $q\upharpoonright(s^{\frown}x)\Vdash\tau=\check{\gamma}_{s}$. Now let $r^{s}:=\{t\in q\upharpoonright s\;|\;|t|\leq|s|\vee t(|s|)\in B_{s}\}$. It follows that $r^{s}\leq_{0}q\upharpoonright(s^{\frown}\alpha)$ and $r^{s}\Vdash\tau=\check{\gamma}_{s}$, i.e. $r^{s}\in D$. Let $r:=\bigcup_{s\in q,|s|=n-1}r^{s}$. Then $r\leq_{0}q$ and whenever $s\in r$, |s|=n1|s|=n-1, $r\upharpoonright s=r^{s}$ and thus $r\upharpoonright s\in D$. This $r$ is what we wanted to construct. ∎
As is common with Prikry-type forcings, $\mathbb{L}(\mu,(\kappa_{n},I_{n},B_{n})_{n\in\omega})$ derives its regularity properties from the Prikry property together with a sufficiently closed direct extension ordering. For technical reasons – most importantly, to show that our later iteration forces instances of $\mathsf{LIP}$ – we need that the direct extension ordering is strongly $\mu-$directed closed. To show this, we first prove some elementary statements regarding ideal combinatorics.
@@
$X\times Y\supseteq A\in I\otimes J:\Leftrightarrow\{x\in X\;|\;\{y\in Y\;|\;(x,y)\in A\}\in J^{+}\}\in I.$
@@
In particular, we may now write expressions such as $I_{0}\otimes\dots\otimes I_{n}$” without using brackets to indicate the order of operations, since e.g. $I_{0}\otimes(I_{1}\otimes(I_{2}\otimes I_{3}))=(I_{0}\otimes I_{1})\otimes(I_{2}\otimes I_{3})$.
The Fubini product is related to Namba forcing as follows: Given finitely many sets $(X_{k})_{k<n}$ and ideals $(J_{k})_{k<n}$, we say that $p\subseteq\bigcup_{l<n}\prod_{k\leq l}X_{k}$ is a $(J_{k})_{k<n}$-Laver-tree if $p$ is nonempty, closed under restriction and for every $s\in p$ with $|s|<n$, $\operatorname{osucc}_{p}(s):=\{x\in X_{|s|}\;|\;s^{\frown}x\in p\}\in J_{|s|}^{+}$. We let $T(p):=\{s\in p\;|\;|s|=n\}$.
@@
Let $J_{0},\dots,J_{n-1}$ be ideals. Then a set $A\subseteq X_{0}\times\dots\times X_{n-1}$ is in $(J_{0}\otimes\dots\otimes J_{n-1})^{+}$ if and only if there is a $(J_{k})_{k<n}$-Laver-tree $p$ such that $T(p)\subseteq A$.